A spring has a force constant of 80 N/m, it vibrates with a frequency of 7 Hz.




Calculate the,




a) period;




b) angular frequency;




c) mass of the spring.

1
Expert's answer
2022-02-01T15:34:01-0500

a)

"T=\frac 1\nu=0.14~s,"

b)

"\omega=2\pi \nu=44~\frac{rad}s,"

c)

"m=\frac k{ 4\pi^2\nu^2}=41.4~g."